Under plan A Giselle would have to pay the first $80 of her medical bills. plus 40% of the rest. Under Plan B Giselle would pay the first $240 but only 25% of the rest. For what amount of medical bills will plan B Save Giselle Money? Assume she has over $240 in bills.
Giselle would save with plan B if she had more then ?? in bills

Answer by mananth(16949)

Under plan A Giselle would have to pay the first $80 of her medical bills. plus 40% of the rest. Under Plan B Giselle would pay the first $240 but only 25% of the rest. For what amount of medical bills will plan B Save Giselle Money? Assume she has over $240 in bills.
...
Let the medical bills be $x
PLAN A
80 of her bill
Balance=
x-80
..
40% 0f (x-80)= 0.4(x-80)
80+0.4(x-80)
...
PLAN B
...
$240 of her bill
Balance = x-240
25% of (x-240)=0.25(x-240)
240+0.25(x-240)
...
80+0.4(x-80)=240+0.25(x-240)
80+0.4x-32=240+0.25x-60
0.4x-0.25x=240-60-80+32
0.15x=132
/0.15
x=132/0.15
x=880
they are same when the bill is $880
However when the bill exceeds $880 PLAN B will save money for her.
..
m.ananth@hotmail.ca